OpenStudy (anonymous):

The equation for a projectile's height versus time is . A tennis ball machine serves a ball vertically into the air from a height of 2 feet, with an initial speed of 130 feet per second. Which equation correctly models the ball's height as a function of time?

OpenStudy (anonymous):

y = y0 + v0 t + (1/2) a t^2 a = -32 ft/s^2 y0 = 2 ft v0 = 130 ft/s